### v2

Abstract
Recently, introducing Tensor Decomposition (TD) techniques into unsupervised feature selection (UFS) has been an emerging research topic. A tensor structure is beneficial for mining the relations between different modes and helps relieve the computation burden. However, while existing methods exploit TD to preserve the data tensor structure, they do not consider the influence of data orientation and thus have difficulty in handling orientation-specific data such as time series. To solve the above problem, we utilize the orientation-dependent tensor-tensor product from Tensor Singular Value Decomposition based on $\star_{\mathbf{M}}$-product (T-SVDM) and extend the one-dimensional Sparse Principal Component Analysis (SPCA) to a tensor form. The proposed sparse tensor PCA model can constrain sparsity at the specified mode and yield sparse tensor principal components, enhancing flexibility and accuracy in learning feature relations. To ensure fast convergence and a flexible description of feature correlation, we develop a convex version specially designed for general UFS tasks and propose an efficient slice-by-slice algorithm that performs dual optimization in the transform domain. Experimental results on real-world datasets demonstrate the effectiveness and remarkable computational efficiency of the proposed method for tensor data of diverse structures over the state-of-the-arts. With a proper combination of data orientation and transform domain, our method is promising for various applications. The codes related to our proposed methods and the experiments are available at https://github.com/zjj20212035/STPCA.git.
keywords:
[…]
1 Introduction
Feature Selection (FS) is a technique of selecting as few features (variables) as possible to meet the demands of a certain task [1, 2]. It is usually employed as a pre-processing algorithm to relieve the computational burden and enhance interpretability by offering only the necessary features. According to the amount of supervised information during training, FS methods can be classified as supervised [3], semi-supervised [4], and unsupervised [5]. Among them, unsupervised feature selection (UFS) is more challenging due to the unavailability of labels. To overcome the difficulty, most UFS methods adopt learning methods including but not limited to principal component analysis (PCA) [38], spectral clustering [40], and graph-based learning [18]. In this paper, we focus on PCA-based UFS methods.
In most UFS methods, data samples are vectorized, resulting in the curse of dimensionality [6] as well as data structure loss when dealing with tensors. As a natural extension of matrix to store multi-way data samples, tensor is now a common processed object in various communities such as image processing [7], signal processing [8], and economical analysis[9]. A $n$-th order tensor comprises $n$ modes, each containing several dimensions. The tensor structure contains rich information on the relations of modes and dimensions, which is beneficial for analyzing the features. In this paper, we focus on leveraging the tensor structure to reduce computational complexity without compromising UFS performance.
To utilize the structure information, tensor-based UFS methods apply all kinds of Tensor Decomposition (TD) techniques[18, 19, 20]. Tensor decomposition aims to factorize a tensor into a product of several subtensors, which has been considerably studied for decades since the famous Tucker decomposition [21] was proposed in 1964. As different definitions for tensor rank are established, various TD methods have been put forward, such as Higher-order Singular Value Decomposition (HOSVD) [12], CANDECOMP/PARAFAC decomposition (CPD) [22, 23], and Tensor Singular Value Decomposition (T-SVD)[24]. One purpose of TD is to mine the relations between modes and dimensions, which are naturally related to UFS. TD methods preserve the tensor structure of UFS, while UFS helps validate the accuracy and interpretability of the relationships learned by the TD method. By minimizing the reconstruction error between a constructed tensor and its decomposition, tensor-based UFS methods successfully preserve the data structure information. However, most of these methods do not develop sparse TD and thus use the classifier matrix to perform UFS, which is not beneficial for interpretability and utilization of the learned element relations. To address this problem, we start by finding the compact representations of the factor matrices. As is widely known, TD methods are also viewed as ways to achieve Tensor Principal Component Analysis (TPCA) [12, 25]. The principal components have the greatest variance among data samples on a global manifold. Their directions describe the feature necessity during the projection onto the manifold. Therefore, the key to selecting interpretable and discriminative features could be hidden in TPCA.
In terms of performing linear projection, the current TPCA methods can be roughly divided into two types: Tuckers-based TPCA [12, 26, 27] and T-SVDM-based TPCA [28, 29, 30]. Tuckers-based TPCA applies direction-unfolding products defined by Tucker decomposition and its variations (including CPD, which can be viewed as a special case of Tucker’s), while T-SVDM-based TPCA is based on the $\star_{\mathbf{M}}$-product defined by T-SVD and its generalization T-SVDM [31]. The main differences between the above two types of methods lie in the following two aspects: 1) Tucker-based methods conduct tensor-matrix products through tensor matricization in the original domain, using factor matrices, while T-SVDM-based methods perform tensor-tensor products in the transform domain with a factor tensor. By applying factor tensors, T-SVDM-based methods obtain significantly more weights assigned for elements in the data tensor, describing feature relations in a more delicate way. 2) Tucker-based methods obtain identical results regardless of data organization, while T-SVDM-based methods can achieve various decompositions by applying different transform domains and changing tensor orientation. The reorientation works as prior knowledge of mode relations and guides the T-SVDM-based methods to further learn the feature relations in a proper direction, especially in some orientation-specific scenarios such as time series. Moreover, the transform operation provides T-SVDM-based methods with greater flexibility, allowing their application to various scenarios. In a word, T-SVDM-based methods have a unique advantage in handling complex scenarios. We are interested in introducing T-SVDM-based TPCA into UFS. To achieve this goal, an extension to a sparse framework is required to ensure the preservation of only the necessary feature relations. To the best of our knowledge, this framework is still absent.
This paper develops a novel sparse T-SVDM-based TPCA framework and deduces a convex version for UFS. We explore a way of leveraging the tensor structure to reduce computational complexity while preserving satisfying UFS performance. Our contributions are summarized as follows:
  1. [•]
  2. 1.
    We put forward a novel Sparse Tensor PCA model based on $\star_{\mathbf{M}}$-Product (STPCA-MP) that extends the principle of one-dimensional Sparse PCA (SPCA) to a tensor form by applying a tensor decomposition technique called T-SVDM. By considering the orientation of the data tensor, the proposed model can better utilize the data organization and learn the feature correlations.
  3. 2.
    We deduce a convex version of STPCA-MP specially designed for unsupervised feature selection tasks by merging the encoder tensor and decoder tensor into the unified sparse reconstruction factor tensor. Our method not only preserves the global information by minimizing the reconstruction error but also demonstrates consistency between TD and UFS by utilizing the reconstruction tensor to perform UFS, fully leveraging learned latent feature relations and gaining more interpretability. We prove that the global optimal solution falls onto the Hermitian Positive Semidefinite Cone.
  4. 3.
    We design an optimization algorithm that exploits the transform domain to solve a dual problem in a slice-by-slice manner. By applying the HPSD projector, we can efficiently solve each subproblem and end up with a score map for feature selection. The convergence of the proposed algorithm is proven theoretically and practically. Compared to non-tensor-based methods, our proposed method reaches a lower computational complexity.
  5. 4.
    Real-world datasets are used to validate the effectiveness and efficiency of the proposed method, especially towards orientation-specific data like time series.
Figure 1: Overall framework of our work
2 Related Work
In this section, we introduce notations used throughout the paper, and research revolving around tensor decomposition, tensor PCA, and UFS.
2.1 Notations
2.1.1 Basic Symbols
In most cases, we use capital letters (e.g. $A$) to denote sets, except the common ones like integer set $\mathbb{Z}$, the real-number set $\mathbb{R}$ and the complex-number set $\mathbb{C}$. In particular, $H^{d}_{+}$ stands for the HPSD Cone for $d\times{d}$ matrices, and $\Phi$ is the empty set. For an integer $n$, $[n]$ represents the set $\left[1,2,3,...,n\right]$. #$\#(A)$ stands for the number of elements in the set $A$. A set with the superscript "o" has ordered elements; for example, $A^{o}=\{1,2,3\}\neq B^{o}=\{1,3,2\}$.
We use lowercase letters (e.g. $a$) to denote scalars, boldface lowercase letters (e.g. $\mathbf{a}$) to denote vectors, boldface capital letters (e.g. $\mathbf{A}$) and Euler script letters (e.g. $\mathcal{A}$) to denote tensors. For a $n^{th}$-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{\cdots}\times{d_{n}}}$, its $(i_{1},i_{2},\cdots,i_{n})$-th element is denoted by $\mathcal{A}_{i_{1}i_{2}\cdots i_{n}}$. For a matrix, the corresponding element is denoted by $a_{i_{1}i_{2}}$. For a matrix $\mathbf{A}$, its row vectors are denoted by $\mathbf{A}_{j,:}$ or boldface lowercase letters $\mathbf{a}^{j}$. The column vectors are denoted by $\mathbf{A}_{:,j}$ or boldface lowercase letters $\mathbf{a}_{j}$.
There are two types of common sections for third-order tensors: fibers and slices. The vector along the $n$-th mode is referred to as an $n$-mode fiber. A third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ has column (1-mode) fibers $\mathcal{A}_{:i_{2}i_{3}}$, row (2-mode) fibers $\mathcal{A}_{i_{1}:i_{3}}$, and tube (3-mode) fibers $\mathcal{A}_{i_{1}i_{2}:}$. A slice is a matrix obtained by fixing all but two indices in a tensor. For a third-order tensor, slices can be divided into frontal slices $\mathcal{A}_{::i_{3}}$, lateral slices $\mathcal{A}_{:i_{2}:}$ and horizontal slices $\mathcal{A}_{i_{1}::}$.
The transpose, conjugation, and conjugate transpose of a matrix are denoted by $\mathbf{A}^{T}$, $\mathbf{A}^{*}$, and $\mathbf{A}^{H}$, respectively. We denote the trace of a square matrix by ${\rm Tr}(\mathbf{A})$. We denote Frobenius norm, nuclear norm and $\ell_{2,1}$ norm by ${\|\mathcal{A}\|}_{F}$,${\|\mathbf{A}\|}_{*}$, and ${\|\mathbf{A}\|}_{2,1}$ respectively. The nuclear norm of a matrix is defined as the sum of its singular values. The $\ell_{2,1}$-norm is obtained by calculating the $\ell_{2}$-norm of each column vector in a matrix and summing the results. The Kronecker product is denoted by $\bigotimes$.
2.1.2 Tensor Operations
In this section, we introduce several tensor operations.
a) Tensor unfolding and tensor-matrix product
We use $n$-mode unfolding to perform tensor matricization. A $m$-th order tensor $\mathcal{A}^{d_{1}\times{d_{2}}\dots\times{d_{m}}}$ can be transformed into $\mathbf{X}_{(n)}$ along the $n$-th mode where
[…]
Using the unfolding technique, the tensor-matrix product is defined as
$\mathcal{Y}=\mathcal{X}\times_{n}\mathbf{U}\iff\mathbf{Y}_{(n)}=\mathbf{U}\mathbf{X}_{(n)},$ (2)
which is also known as the $n$-mode product.
b) $\star_{\mathbf{M}}$-product
In the T-SVDM framework, third-order tensor products are achieved through the $\star_{\mathbf{M}}$-product. Given a third-order tensor $\mathcal{A}\in\mathbb{C}^{d\times{m}\times{n}}$, we first denote a tensor in the transform domain specified by an invertible matrix $\mathbf{M}\in\mathbb{C}^{n\times{n}}$:
$\hat{\mathcal{A}}:=\mathcal{A}\times_{3}\mathbf{M}.$ (3)
Then, given another tensor $\mathcal{B}\in\mathbb{C}^{m\times{l}\times{n}}$, $\mathcal{A}\star_{\mathbf{M}}\mathcal{B}=\mathcal{C}$ is defined as
$\hat{\mathcal{C}}_{::i}$ $=\hat{\mathcal{A}}_{::i}\hat{\mathcal{B}}_{::i},\quad i=1,2,\cdots,n,$ (4)
$\mathcal{C}$ $=\hat{\mathcal{C}}\times_{3}\mathbf{M}^{-1}\in{\mathbb{C}^{d\times{l}\times{n}}}.$
Hereafter, we use $\hat{\mathcal{A}}$ to denote the result of $\mathcal{A}\times_{3}\mathbf{M}$. The $\star_{\mathbf{M}}$-product is a slice-by-slice operation and is thus orientation-dependent. Different orientations of the tensor will lead to different results.
The identity tensor $\mathcal{I}$ under $\star_{\mathbf{M}}$ satisfies the following condition:
$\mathcal{A}\star_{\mathbf{M}}\mathcal{I}=\mathcal{I}\star_{\mathbf{M}}\mathcal{A}^{H}=\mathcal{A}.$ (5)
Each frontal slice of $\hat{\mathcal{I}}$ is an identity matrix $\mathbf{I}$.
The conjugate transpose of $\mathcal{A}\in\mathbb{C}^{d\times{m}\times{n}}$ under $\star_{\mathbf{M}}$ is defined as:
$\left(\hat{\mathcal{A}}^{H}\right)_{::i}=\left(\hat{\mathcal{A}}_{::i}\right)^{H},\quad i=1,\cdots,n,$ (6)
which makes sure that $\mathcal{A}^{H}\star_{\mathbf{M}}\mathcal{B}^{H}=\left(\mathcal{B}\star_{\mathbf{M}}\mathcal{A}\right)^{H}$. We can further define a $\star_{\mathbf{M}}$-orthogonal tensor $\mathcal{A}\in{\mathbb{C}^{m\times{m}\times{n}}}$ if
$\mathcal{A}^{H}\star_{\mathbf{M}}\mathcal{A}=\mathcal{A}\star_{\mathbf{M}}\mathcal{A}^{H}=\mathcal{I}.$ (7)
2.2 T-SVDM
Using the $\star_{\mathbf{M}}$ product, T-SVDM [31] decomposes a third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ into:
$\mathcal{A}=\mathcal{U}\star_{\mathbf{M}}\mathcal{S}\star_{\mathbf{M}}\mathcal{V}^{H},$ (8)
where $\mathcal{U}\in\mathbb{C}^{d_{1}\times d_{1}\times d_{3}}$, $\mathcal{V}\in\mathbb{C}^{d_{2}\times d_{2}\times d_{3}}$ are $\star_{\mathbf{M}}-unitary$, and $\mathcal{S}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ is a f-diagonal tensor whose frontal slices are diagonal. The decomposition is called T-SVD [24] when $\mathbf{M}$ is the discrete Fourier transform (DFT) matrix and T-SVDM [31] in a usual case.
[28] was the first to establish the relationship between T-SVD and tensor PCA. It applies T-SVD to compress the human-face images with a pre-rotation of the data tensor. [31] further extends this idea to T-SVDM. [29] highlights the importance of leveraging the orientation-dependence of T-SVDs to extract information from different directions. Inspired by these works, we derive a sparse model for tensors combining T-SVDM and SPCA.
2.3 PCA and Sparse PCA (SPCA)
Principal Component Analysis (PCA) [47] is a well-known unsupervised feature extraction method. It aims to find an orthogonal linear transformation matrix that projects the original data vectors into a lower-dimensional manifold, maximizing the variance of the projected samples. The optimization problem of PCA can be formulated as the following self-contained regression form:
$\min\limits_{\mathbf{U}^{T}\mathbf{U}=\mathbf{I}_{k}}{\|{\mathbf{X}-\mathbf{U}\mathbf{U}^{T}\mathbf{X}}\|}^{2}_{F},$ (9)
where $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\cdots,\mathbf{x}_{n}\right]\in\mathbb{R}^{d\times{n}}$ is the data matrix, with each column representing a sample. $\mathbf{U}\in\mathbb{R}^{d\times{k}}$ ($k\ll d$) is an orthogonal transformation matrix.
To obtain sparse weights and improve interpretability, Sparse PCA (SPCA) [32] was proposed by adding ridge and lasso constraints to PCA:
$\min\limits_{\mathbf{U},\mathbf{Q}}$ ${\|{\mathbf{X}-\mathbf{U}\mathbf{Q}^{T}\mathbf{X}}\|}^{2}_{F}+\alpha\sum\limits_{j=1}^{k}{\|{\mathbf{q}_{j}}\|}^{2}_{2}+\beta\sum\limits_{j=1}^{k}{\|{\mathbf{q}_{j}}\|}_{1}$ (10)
$s.t.$ ${\mathbf{U}^{T}\mathbf{U}=\mathbf{I}_{k}},$
where $\alpha,\beta>0$ are regularization parameters.
The self-contained regression term in SPCA incorporates an encoder-decoder architecture, selectively transforming a sparse subset of input features into encoded latent representations and reconstructing them through an orthogonal decoder, as Figure 2 shows. This approach gains flexibility by relaxing the transformation matrix $\mathbf{Q}$. When $\mathbf{Q}=\mathbf{U}$, SPCA reduces to PCA.
Figure 2: The encoder-decoder principle of one-dimensional SPCA
By minimizing reconstruction errors, SPCA preserves the global data structure in an unsupervised manner, maintaining essential geometric relationships within the data manifold. The lasso constraint enables SPCA to assign weights only to the necessary features. The orthogonality constraint ensures that the latent features are complementary and reduces redundancy. Therefore, SPCA is well-suited for UFS tasks, particularly in balancing the clustering performance and feature interpretability.
2.4 Tensor-based unsupervised feature selection
Tensor-based unsupervised feature selection remains an emerging research area. Most of the existing methods rely on domain knowledge and are limited to certain application scenarios. For example, [20] proposes an image-transformation-based feature selection method, aiming at vehicle detection. [34] constructs a multi-omics data tensor and exploits feature selection criterion designed using the knowledge in bioinformatics. Recently, some methods designed for general cases have emerged. GRLTR [35] combines graph regularization with tensor tubal rank constraint to simultaneously preserve global and local information. However, it still requires vectorization. CPUFS [18] derives a model based on CPD and effectively preserves the tensor structure by minimizing the reconstruction error. It represents a promising approach to tensor-based UFS. However, CPUFS does not derive sparse factor matrices. Instead, it applies spectral clustering and utilizes a sparse linear classifier matrix to perform UFS, leading to insufficient utilization of the information in factor matrices and reduced interpretability of selected features. Additionally, the constructed graph in CPUFS needs the number of classes as prior knowledge and can be unreliable due to noise features.
Recently, developing sparse PCA models for UFS appears to be a new trend [36, 37, 38, 39]. PCA-like methods do not rely on the construction of graphs. By global manifold learning, they can preserve the data structure and achieve superior average performance. What’s more, they utilize the factor matrix that maximizes the variance of data samples in the projected space, whose sparse weights demonstrate a clearer relevance of the original features and have better interpretability. Unfortunately, existing PCA-like methods are non-tensor-based. Motivated by the above considerations, this study develops a sparse TPCA-based methodology tailored for UFS tasks.
3 Methodology
In this section, we extend the framework of SPCA to tensor form by utilizing techniques from T-SVDM and derive a model called Sparse Tensor PCA based on $\star_{\mathbf{M}}$ product (STPCA-MP). Then, we deduce a convex version of STPCA-MP for UFS tasks by merging the encoder and decoder tensors into a single sparse reconstruction tensor. We propose an efficient slice-by-slice optimization algorithm that individually solves each dual subproblem in the transform domain. Eventually, by considering both the transform matrix $\mathbf{M}$ and the learned reconstruction tensor, we compute a feature score map and perform feature selection. The overall framework is illustrated in Figure 1.
3.1 Model formulation of STPCA-MP
We formulate the model in the third-order tensor scenario for the simplicity of deduction and algorithm design. The principle can be extended to high-order scenarios in a recursive fashion. See Section 4 for details.
Given a data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ containing $n$ samples of size $d_{1}\times d_{2}$, we can reorient it with an order set $D^{o}$ such that the second mode represents the sample dimension. We denote the rotated data tensor by $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$. Based on (8), the row dimensionality of each frontal slice in $\mathcal{X}_{D^{o}}$ can be reduced with $\mathcal{U}^{H}\in\mathbb{C}^{k\times q\times p}(k<d)$, followed by reconstruction with $\mathcal{U}$:
$\mathcal{Y}$ $=\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}},$ (11)
$\tilde{\mathcal{X}}_{D^{o}}$ $=\mathcal{U}\star_{\mathbf{M}}\mathcal{Y},$
where $\mathcal{Y}\in\mathbb{C}^{k\times n\times p}$, and $\tilde{\mathcal{X}}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$ is the reconstructed tensor.
Intuitively, we can first formulate a tensor PCA model following the idea of one-dimensional PCA in (9):
$\underset{\mathcal{U}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (12)
$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
Then, inspired by the encoder-decoder structure of SPCA, we use another tensor $\mathcal{Q}\in\mathbb{C}^{d\times k\times p}$ to perform feature extraction instead of $\mathcal{U}$. By doing so, we obtain a more flexible reconstruction framework that yields a better latent feature representation:
$\underset{\mathcal{U}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (13)
$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
Specifically, $\mathcal{Q}$ contains the weights that measure the importance of each original feature during the projection into the latent feature space. To encode only the necessary original features, $\mathcal{Q}$ is constrained to have sparse loadings. Meanwhile, for highly relevant features, we hope the model can treat them equally, instead of preferring certain features, especially when $dm\gg n$. Therefore, lasso regression and ridge regression terms are introduced as constraints for $\mathcal{Q}$, resulting in a sparse Tensor PCA model:
$\underset{\mathcal{U},\mathcal{Q}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}+\alpha\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|^{2}_{2}}+\beta\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|_{1}}$ (14)
$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
By obtaining the optimal $\mathcal{Q}$, the sparse tensor principal components can be extracted from the data tensor.
Discussion: To this end, the proposed model is directly applicable to tasks traditionally addressed by SPCA (e.g., image compression, visualization, and denoising) while preserving the tensor structure. However, the encoder tensor alone is suboptimal for UFS tasks. In the following section, we reformulate the model to meet UFS-specific requirements.
3.2 Convex version of STPCA-MP for UFS
As previously discussed, the self-contained regression term has an encoder-decoder structure. Therefore, the importance of each original feature is not only expressed by the weights contained in $\mathcal{Q}$ (encoder) but also described by $\mathcal{U}$ (decoder). Comprehensively considering their relationship is essential for adapting the model to UFS.
Dual Problem: Since the transformation between tensor $\mathcal{A}$ and its counterpart $\hat{\mathcal{A}}$ is uniquely defined by $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}$, Problem (14) has its dual problem in the transform domain, which means they can be solved simultaneously. To facilitate optimization, we can rewrite Problem (14) in the transform domain where matrix algebra can be applied:
$\underset{\hat{\mathcal{U}},\hat{\mathcal{Q}}}{\min}$ $\sum\limits^{p}_{i=1}{\|\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}-\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\|^{2}_{F}}+\lambda\sum\limits^{p}_{j=1}{\|\hat{\mathcal{Q}}_{::i}\|_{2,1}}$ (15)
$s.t.\quad\hat{\mathcal{U}}_{::i}^{H}\left(\hat{\mathcal{U}}_{::i}\right)=\mathbf{I},i\in[p].$
The $\ell_{2,1}$ norm is employed to simultaneously achieve the goals of lasso regression and ridge regression.
Variable Merging: Notably, there exists a tensor $\mathcal{A}\in\mathbb{C}^{d\times d}$ whose $i$-th frontal slice in the transform domain $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$. Each row vector in $\hat{\mathcal{A}}_{::i}$ contains reconstruction weights that combine elements from both $\hat{\mathcal{U}}_{::i}$ and $\hat{\mathcal{Q}}^{H}_{::i}$. Therefore, $\mathcal{A}$ is a more suitable feature selection tensor than $\mathcal{Q}$. Guided by $D^{o}$, each frontal slice in $\hat{\mathcal{A}}$ reconstructs the fibers of a certain mode in the data tensor, utilizing the structural information hidden in the mode organization. Remarkably, it can be proven that the optimal $\hat{\mathcal{A}}$ is a Hermitian positive semidefinite (HPSD) matrix, detailed by Theorem 1.
Theorem 1.
Given the optimal solution $\hat{\mathcal{U}}^{opt}_{::i}$ and $\hat{\mathcal{Q}}^{opt}_{::i}$ to Problem (15), then $\hat{\mathcal{A}}^{opt}_{::i}=\hat{\mathcal{U}}^{opt}_{::i}(\hat{\mathcal{Q}}^{opt}_{::i})^{H}\in H^{q}_{+}$.
[…]
pfProof of Theorem 1 {pf} See Appendix.
Norm Simplification: Since $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$ and $d\gg k$, $\hat{\mathcal{A}}_{::i}$ is also a low-rank matrix. It is widely known that nuclear norm $\|\cdot\|_{*}$ is often used as the convex approximation to the low-rank matrix constraint. However, its high computational complexity poses a challenge. Fortunately, in our case, $\|\hat{\mathcal{A}}_{::i}\|_{*}={\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ due to the HPSD constraint, significantly reducing computational complexity.
Meanwhile, provided that $\hat{\mathcal{U}}_{::i}$ is orthogonal, it is easy to prove that
${\|\hat{\mathcal{Q}}_{::i}\|}_{2,1}={\|\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}\|}_{2,1}={\|\hat{\mathcal{A}}_{::i}\|}_{2,1}={\|\hat{\mathcal{A}}^{H}_{::i}\|}_{2,1}.$ (16)
Therefore we can use ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ to simplify optimization. Moreover, minimizing ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ is equivalent to minimizing ${\|\mathcal{A}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\mathcal{A}_{::i}\right)$ when $\mathbf{M}$ is orthogonal. This implies that the elements in a column vector of $\mathcal{A}_{::i}$ (representing different weights for the same feature) are minimized toward zero for redundant features, thereby fulfilling the feature selection task.
Finally, a convex version of STPCA-MP for UFS is formulated as follows:
$\underset{\mathcal{A}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{A}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}+\lambda\sum\limits^{p}_{i=1}{\|\mathcal{A}}_{::i}\|_{2,1}+\eta\sum\limits^{p}_{i=1}{\rm Tr}\left(\mathcal{A}_{::i}\right)$ (17)
$s.t.$ $\hat{\mathcal{A}}_{::i}\in H^{q}_{+},i\in[p].$
Discussion: The convex version of STPCA-MP exhibits a compact model expression. By merging the sparse encoder and orthogonal decoder, the reconstruction tensor simultaneously assigns weights to original features and their latent representations, comprehensively measuring feature importance.
3.3 Optimization Algorithm
The key to solving Problem (17) lies in leveraging the tensor structure to handle the HPSD constraint with reduced computational complexity.
Multiple Dual Subproblems: The proposed model in the original domain (Problem (17)) has a dual problem in the transform domain, since their solutions are uniquely connected via $\mathbf{M}$. As demonstrated below, the model is significantly easier to solve in the transform domain. First, we rewrite Problem (17) in the transform domain:
$\underset{\hat{\mathcal{A}}}{\min}$ $\sum\limits^{p}_{i=1}\left({\|\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}-\hat{\mathcal{A}}_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\|^{2}_{F}}+\lambda{\|\hat{\mathcal{A}}_{::i}\|}_{2,1}+\eta{\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)\right)$ (18)
$s.t.$ $\hat{\mathcal{A}}_{::i}\in H^{q}_{+},i\in[p],$
which can be decomposed into $p$ independent subproblems. We can divide and conquer each subproblem and gather the results to construct $\mathcal{A}$.
Derivative: For the $i$-th subproblem, we calculate the derivative concerning $\hat{\mathcal{A}}_{::i}$ and set the value to 0, yielding:
$\hat{\mathcal{A}}_{::i}=\left(\mathbf{S}_{i}-\frac{\eta_{i}}{2}\mathbf{I}_{q}\right)\left(\mathbf{S}_{i}+\lambda_{i}{\mathbf{W}_{i}}+\epsilon_{2}\mathbf{I}_{q}\right)^{-1},$ (19)
where $\mathbf{S_{i}}=\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)^{T}_{::i}$ and $\mathbf{W}_{i}\in{\mathbb{R}^{q\times q}}$ is a diagonal matrix whose $j$-th diagonal element is calculated by:
$\mathbf{w}_{jj}=\left(1/\left(2\sqrt{\hat{\mathbf{a}}_{j}^{H}\hat{\mathbf{a}}_{j}+\epsilon_{1}}\right)\right).$ (20)
HPSD Projection: The next step is to project $\hat{\mathcal{A}}_{::i}$ onto the HPSD cone. This is achieved using the HPSD projection operator $P_{H^{d}_{+}}(\cdot)$ [55] which, for an arbitrary square matrix $\mathbf{A}\in\mathbb{C}^{b\times b}$, is defined as:
$P_{H^{b}_{+}}(\mathbf{A})$ $=\Pi_{H^{b}_{+}}\left(\Pi_{H^{b}}(\mathbf{A})\right)$ (21)
$=\Pi_{H^{b}_{+}}\left(\frac{1}{2}(\mathbf{A}+\mathbf{A}^{H})\right)$
$=\sum\limits_{i=1}^{b}max\left\{\sigma_{i},0\right\}\mathbf{u}_{i}^{H}\mathbf{u}_{i},$
where $\sum\nolimits_{i=1}^{b}\sigma_{i}\mathbf{u}_{i}^{H}\mathbf{u}_{i}$ is the eigenvalue decomposition of $\frac{1}{2}(\mathbf{A}+\mathbf{A}^{H})$.
Based on the above techniques, we solve the $i$-th subproblem of (18) with the following two-stage algorithm:
Step 1: Fix $\mathbf{W}_{i}$ and Update $\hat{\mathcal{A}}_{::i}$.
When $\mathbf{W}_{i}$ is fixed, the optimal $\hat{\mathcal{A}}_{::i}$ is obtained at:
$\hat{\mathcal{A}}_{::i}=P_{H^{q}_{+}}\left(\left(\mathbf{S}-\frac{\eta_{i}}{2}\mathbf{I}_{q}\right)\left(\mathbf{S}+\lambda_{i}{\mathbf{W}_{i}}+\epsilon_{2}\mathbf{I}_{q}\right)^{-1}\right).$ (22)
Step 2: Fix $\hat{\mathcal{A}}_{::i}$ and Update $\mathbf{W}_{i}$.
When $\hat{\mathcal{A}}_{::i}$ is fixed, $\mathbf{W}_{i}$ can computed directly by (20).
Discussion: Formulating a series of convex subproblems simplifies the optimization algorithm design. Although the HPSD projection requires eigenvalue decomposition with a computational complexity of $O(x^{3})$, the exploitation of tensor structures ensures that $x\ll\#{\rm Feature}$. The efficiency of the proposed method is demonstrated in the experimental results.
3.4 Feature selection
We assess the importance of an original feature by its contribution to the reconstruction. Given the presence of the transform domain, the influence of the transform matrix $\mathbf{M}$ must be considered.
Suppose we reorient the data tensor to obtain $\mathcal{X}_{D^{o}}\in\mathbb{C}^{p\times n\times q}$. For the $ij$-th feature $f_{ij}$, its transform-domain $\hat{f}_{ij}$ counterpart and reconstructed version $\tilde{f}_{ij}$ are expressed as:
$\hat{f}_{ij}$ $=\sum\limits^{q}_{h=1}m_{jh}f_{ih},\quad\text{and}$ (23)
$\tilde{f}_{ij}$ $=\sum\limits^{p}_{l=1}\hat{a}_{ilj}\hat{f}_{lj}=\sum\limits^{p}_{l=1}\left(\sum\limits^{p}_{k=1}m_{ik}a_{ilk}\right)\left(\sum\limits^{q}_{h=1}m_{jh}f_{lh}\right)=\sum\limits^{p}_{l=1}\sum\limits^{q}_{h=1}\left(\sum\limits^{p}_{k=1}m_{jh}m_{ik}a_{ilj}\right)f_{lh},$
which indicates that the contribution of $f_{lh}$ to the reconstruction of $f_{ij}$ can be described by $\sum^{p}_{k=1}m_{jh}m_{ik}a_{ilj}$. Therefore, the contribution matrix of $f_{lh}$ is defined as:
$\left[Contribution\right]_{ij}=\sum\limits^{p}_{l=1}\sum\limits^{q}_{h=1}\left(\sum\limits^{p}_{k=1}m_{jh}m_{ik}a_{ilj}\right).$ (24)
Finally, the score for $f_{lh}$ is computed as $\|Contribution\|^{2}_{F}$, which comprehensively captures its influence on the reconstruction of all features and leverages the global information. We summarize the whole algorithm in Algorithm 1.
Discussion: Equation (24) shows that an appropriate choice of $\mathbf{M}$ integrates feature correlation information from other modes into the mode targeted for slice-by-slice optimization. In our proposed method, each feature is scored by globally modeling its dependencies with other features without requiring sample vectorization, thus preserving the data organization and reducing computational complexity.
Algorithm 1 The optimization algorithm for STPCA-MP
0:  A third-order centralized data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$, the order set $D^{o}$ determining $q$, an invertible matrices $\mathbf{M}\in\mathbb{C}^{q\times q}$, regularization parameters $\lambda,\eta$.
0:  $h$ selected features.
1:  Randomly initialize $\left(\tilde{\mathcal{A}}\right)_{::i}\in{H^{q}_{+}}$, transform $\mathcal{X}$ into $\tilde{\mathcal{X}}$.
2:  for $i=1,2,\cdots,q$ do
3:     revolve $\tilde{\mathcal{X}}$ with $D^{o}$ and obtain $\tilde{\mathcal{X}}_{D^{o}}$.
4:     input $\left(\tilde{\mathcal{X}}_{D^{o}}\right)_{::i}$ and alternatively update $\hat{\mathcal{A}}_{::i}$ and $\mathbf{W}_{i}$ according to (22) until converge.
5:  end for
6:  Compute $\mathcal{A}=\hat{\mathcal{A}}\times_{3}\mathbf{M}^{-1}$ and $\|Contribution\|^{2}_{F}$ for each score according to (24).
7:  Output a feature score map $Score$.
8:  Sort $Score$ in descending order, and select features of the $h$ greatest scores. For slice-wise tensor, the score for each dimension will be obtained by summation.
4 Discussion
In this section, we discuss the convergence, computational complexity, and extension to higher-order tensor data of STPCA-MP.
4.1 Convergence analysis
First, we prove the convergence of the proposed algorithm. Let us introduce a lemma [40]:
Lemma 1.
For any nonzero vectors $\mathbf{a}$, $\mathbf{b}\in{\mathbb{R}^{c\times{1}}}$, the following inequality holds:
${\|\mathbf{a}\|_{2}}-\frac{{\|\mathbf{a}\|_{2}^{2}}}{2{\|\mathbf{b}\|_{2}}}\leq{\|\mathbf{b}\|_{2}}-\frac{{\|\mathbf{b}\|_{2}^{2}}}{2{\|\mathbf{b}\|_{2}}}.$ (25)
Based on Lemma 1, we have the following theorem:
Theorem 2.
the objective function value of each subproblem in Problem (18) is non-increasing in each iteration of Algorithm 1 and can converge to a global minimum.
[…]
pf2Proof of 2 {pf2} For simplicity, we temporally use $\mathbf{A}$ and $\mathbf{X}$ to replace $\hat{\mathcal{A}}_{::i}$ and $\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}$. Denote the updated $\mathbf{A}$ as $\mathbf{\tilde{A}}$ and fix $\mathbf{W}$. By rewriting the $i$-th subproblem in (18) with the trace function, we have the following inequality
${\rm Tr}\left(\mathbf{\tilde{A}}\mathbf{S}\mathbf{\tilde{A}}^{H}\right)$ $-2Tr\left(\mathbf{S}\mathbf{\tilde{A}}\right)+\lambda{\rm Tr}(\mathbf{\tilde{A}}\mathbf{W}\mathbf{\tilde{A}}^{H})+\eta{\rm Tr}(\mathbf{\tilde{A}})$ (26)
$\leq{\rm Tr}(\mathbf{A}\mathbf{S}\mathbf{A}^{H})-2Tr(\mathbf{S}\mathbf{A})+\lambda{\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H})+\eta{\rm Tr}(\mathbf{A}).$
Denote $J(\mathbf{A})={\rm Tr}(\mathbf{A}\mathbf{S}\mathbf{A}^{H})-2Tr(\mathbf{S}\mathbf{A})+\lambda{\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H})+\eta{\rm Tr}(\mathbf{A})$ and $J(\mathbf{\tilde{A}})$ as the updated $J(\mathbf{A})$. Inequality (26) can be rewritten as
$J(\mathbf{\tilde{A}})+\lambda{\rm Tr}(\mathbf{\tilde{A}}\mathbf{W}\mathbf{\tilde{A}}^{H})\leq J(\mathbf{A})+\lambda{\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H}).$ (27)
Adding the same item $\sum\limits_{j=1}^{d}{\lambda\epsilon}/{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}$ to both sides of (27), we obtain
$J(\mathbf{\tilde{A}})$ $+\lambda{\rm Tr}(\mathbf{\tilde{A}}\mathbf{W}\mathbf{\tilde{A}}^{H})+\sum\limits_{j=1}^{d}\frac{{\lambda\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}\leq J(\mathbf{A})+\lambda{\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H})+\sum\limits_{j=1}^{d}\frac{{\lambda\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (28)
In the meantime, we have
${\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H})={\rm Tr}(\mathbf{A}^{H}\mathbf{W}\mathbf{A})=\sum\limits_{j=1}^{d}\frac{{\lambda{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (29)
The same equality holds for ${\rm Tr}(\mathbf{\tilde{A}}\mathbf{W}\mathbf{\tilde{A}}^{H})$. We substituted (29) into (28) and get
$J(\mathbf{\tilde{A}})+\sum\limits_{j=1}^{d}\frac{{\lambda({\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon)}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}\leq J(\mathbf{A})+\sum\limits_{j=1}^{d}\frac{{\lambda({\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon)}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (30)
According to Lemma 1, we have
$\sqrt{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}-\frac{{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}\leq\sqrt{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}-\frac{{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (31)
Further, we can accumulate (30) from $j=1$ to $j=d$. Considering $\lambda>0$, we can get
$\lambda\sum\limits_{j=1}^{d}\sqrt{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}-\lambda\sum\limits_{j=1}^{d}\frac{{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}\leq\lambda\sum\limits_{j=1}^{d}\sqrt{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}-\lambda\sum\limits_{j=1}^{d}\frac{{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (32)
By summing (30) and (32), we have
$J(\mathbf{\tilde{A}})+\lambda\sum\limits_{j=1}^{d}\sqrt{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}\leq J(\mathbf{A})+\lambda\sum\limits_{j=1}^{d}\sqrt{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}.$ (33)
Finally, we can get the following inequality
$J(\mathbf{\tilde{A}})+\lambda{\|{\mathbf{\tilde{A}}}\|_{2,1}}\leq J(\mathbf{A})+\lambda{\|{\mathbf{A}}\|_{2,1}}.$ (34)
Each subproblem in (18) is convex and bounded below by 00, ensuring convergence to the global minimum. Since the subproblems are solved independently, Problem (18) is guaranteed to converge to a global solution. When the solution in the transform domain is obtained, the dual problem in the original domain is solved as well.
4.2 Computational complexity
Suppose $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ is rotated into $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$ ($n$ is the number of samples, $(p,q)\in\{(d_{1},d_{2}),(d_{2},d_{1})\}$). For each subproblem, it takes $O\left(q^{2}\right)$ to convert data into the transform domain and initialize the data covariance matrix, and $O\left(q^{3}\right)$ to update the reconstruction matrix in each iteration (including matrix inversion and HPSD projection). Therefore we have $p$ subproblems each solved at the computational cost of $O\left(q^{3}t_{i}+q^{2}n\right)$ ($t_{i}$ is the number of iterations for the $i$-th subproblem). The final feature score computation takes $O(p^{2}q)$. Eventually, the total computational complexity is $O\left(\sum\limits^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)$.
The number of features ($q\times p$) is usually much greater than the number of dimensions in a single mode ($q$ or $p$). Due to the exploitation of tensor structures with separate operations on each slice, STPCA-MP can handle data tensors at a remarkably smaller computation cost than non-tensor-based methods. For example, for image data of size $32\times 32$, non-tensor-based methods requiring matrix inversion or eigenvalue decomposition (such as SOGFS [40] and SPCAFS [38]; see Table 1) have a computational complexity of $O(1024^{3})$, whereas STPCA-MP requires only $O(32\times(32^{3}))$, reducing the complexity by $1024$ times. In addition, the computational complexity is linear to the number of samples, which means the proposed method will still be efficient when facing data tensors with numerous samples.
Table 1 shows the computational complexity of competitive methods in the following experiments. We can see that tensor-based methods reduce the computational burden from $\prod_{i}d_{i}^{x}$ to $d_{i}^{x}$ in each iteration by utilizing the tensor structure. Among tensor-based methods, CPUFS and CPUFSnn are the most recently proposed, which have a computational complexity linear to $d_{i}$ at the price of squared to $n$. They are non-convex models and usually take hundreds of iterations to converge, making it a challenge when facing large-scale datasets.
Table 1: Computational complexity of competitive methods on third-order tensors. $d=d_{1}d_{2}$ is the number of features. $t_{i}$ stands for the number of iterations. $c$ stands for the number of classes.
Method Input Computational complexity
SOGFS [40] matrix $O\left(dn^{2}+t(d^{3}t_{1}+n^{2}c)\right)$
RNE [50] matrix $O\left(dn^{2}+t(t_{1}(n+s)d^{2}+t_{2}n))\right)$
SPCAFS [39] matrix $O\left(d^{2}n+d^{3}t\right)$
NLGMS [56] matrix $O\left(t\left(d^{3}+n^{3}+nd+n^{2}d+d\log{d}\right)\right)$
CPUFS [18] tensor $O\left(d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+n^{2}c)t\right)$
CPUFSnn [18] tensor $O\left((d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+(n_{1}c)^{2}+(n_{2}c)^{2}+n^{2}c)t\right)$
MSPCA [33] tensor $O\left(t_{2}\sum_{i=1}^{2}(t_{1}+n)d_{i}^{3}\right)$
STPCA-MP tensor $O\left(\sum^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)(q\in\{d_{1},d_{2}\})$
4.3 Extension to higher-order tensor data
By conducting the $\star_{\mathbf{M}}$-product in a recursive fashion, the proposed method can be extended to higher-order tensors. Given two $n$-th order tensors $\mathcal{A}\in\mathbb{C}^{d_{1}\times d_{2}\cdots\times d_{n-1}\times d_{n}}$ and $\mathcal{B}\in\mathbb{C}^{d_{2}\times h\cdots\times d_{n}}$, along with a series of invertible matrices $\mathbf{M}_{i}\in\mathbb{C}^{d_{i+2}\times d_{i+2}}$, the $\star_{\mathbf{M}}$-product for higher-order tensors is defined as shown in Algorithm 2. Using this technique, STPCA-MP can be formulated for higher-order settings. However, sequential $\star_{M}$-products with different orientations introduce significant optimization. This remains an open problem under active investigation. Therefore, this paper focuses on the third-order case.
Algorithm 2 $\star_{\mathbf{M}}$-product for higher-order tensor data
0:  $\mathcal{A}\in\mathbb{C}^{d_{1}\times d_{2}\cdots\times d_{n-1}\times d_{n}}$, $\mathcal{B}\in\mathbb{C}^{d_{2}\times h\cdots\times d_{n}}$, $\mathbf{M}_{i}\in\mathbb{C}^{d_{i+2}\times d_{i+2}}$
0:  $\mathcal{C}=\mathcal{A}\star_{M}\mathcal{B}$
1:  Project the input data into the transform domain: $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}_{1}\times_{4}\mathbf{M}_{2}\cdots\times_{n}\mathbf{M}_{n-2}$, ^=×$\hat{\mathcal{B}}=\mathcal{B}\times_{3}\mathbf{M}_{1}\times_{4}\mathbf{M}_{2}\cdots\times_{n}\mathbf{M}_{n-2}$
2:  for $i=1,2,\cdots,\prod\limits^{n}_{j=3}d_{j}$ do
3:     $\hat{\mathcal{C}}=\hat{\mathcal{A}}_{::i_{3}\cdots i_{n}}\hat{\mathcal{B}}_{::i_{3}\cdots i_{n}}$
4:  end for
5:  Project the result back to the original domain $\mathcal{C}=\hat{\mathcal{C}}\times_{n}\mathbf{M}^{-1}_{n-2}\times_{n-1}\mathbf{M}^{-1}_{n-1}\cdots\times_{3}\mathbf{M}^{-1}_{1}$.
6:  Output $\mathcal{C}$
5 Experiments
In this section, we conduct experiments on real-world data to demonstrate our proposed method’s effectiveness. The codes related to our proposed methods and the experiments are available at https://github.com/zjj20212035/STPCA.git.
5.1 Experimental Settings
5.1.1 Experimental environment
All the experiments are conducted on a personal computer with a 2.20 GHz CPU and 64GB main memory under the environment of the Windows 11 operation system.
5.1.2 Comparative methods
We compare our proposed methods with some state-of-the-art UFS methods as follows:
  1. [•]
  2. 1.
    CPUFS and CPUFSnn [18] (2023): CPUFS is a tensor-based UFS method combining CP decomposition and spectral clustering. It utilizes a sparse classifier to perform UFS. CPUFSnn is a variation with nonnegative constraints on the classifier matrix.
  3. 2.
    MSPCA [33] (2014): MSPCA is a sparse tensor PCA method that utilizes Tucker decomposition to make use of tensor structure. It was originally proposed as a feature extraction method. We use the Kronecker product of its factor matrices to perform UFS, as an example of Tucker-based UFS methods.
  4. 3.
    NLGMS [56] (2024): NLGMS is a non-tensor-based method that introduces a self-representation technique into the graph-based framework to preserve global and local data structure. The self-representation term reconstructs each sample with the linear combination of other samples.
  5. 4.
    SPCAFS [38] (2021): SPCAFS is a non-tensor-based method that introduces $\ell_{2,p}$-norm into the one-dimensional PCA model. It forms a simple yet effective model.
  6. 5.
    RNE [50] (2020): RNE is a non-tensor-based method that makes use of the locally linear embedding algorithm to obtain an optimal projection matrix. It keeps the samples close to their neighbors after being projected onto a local manifold.
  7. 6.
    SOGFS [40] (2019): SOGFS is a non-tensor-based method that optimizes an adaptive graph to preserve the local geometrical structure of data. It updates the similarity matrix during each iteration to obtain a more flexible graph.
We also perform clustering with all features as a baseline method. All the comparative methods are implemented using the codes provided by their authors.
5.1.3 Parameter settings
For all methods, we tune the regularization parameters via grid search over $\{10^{-2},10^{-1},1,10,10^{2}\}$. For graph-based methods, we set the number of k-nearest neighbors to 5. For methods utilizing $\ell_{2,p}$-norm, we set $p=1$. For STPCA-MP, we set $\mathbf{M}=\mathbf{I}$ in most experiments except the one that discusses the choice of $\mathbf{M}$. The direction set $D^{o}$ is also tuned by a grid-search strategy in the range $\{\{1,3,2\},\{2,3,1\}\}$. The other parameters remain the default in the open code provided by the authors. The best results of all methods are presented.
5.1.4 Evaluation methodology and metrics
In clustering experiments, after a UFS method has finished selecting features, we evaluate its performance by first conducting K-means clustering and then mapping the obtained pseudo labels to the real labels by adopting the Kuhn-Munkres algorithm [41]. Due to the K-means’ dependence on initialization, we repeat the clustering 30 times. The clustering metrics we use are as follows:
  1. [•]
  2. 1.
    Accuracy (ACC): ACC refers to the clustering accuracy and is computed according to (35) and (36), where $n$ is the total number of samples, $map(i)$ denotes the clustering label of the $i$-th sample after mapping to the real label, and $label(i)$ represents the ground truth label. A larger ACC indicates better performance.
    $\text{ACC}=\frac{1}{n}\sum\limits_{i=1}^{n}\delta\left(map(i),label(i)\right).$ (35)
    $\delta(p,q)=\left\{\begin{aligned} &1\quad\text{if}\;p=q,\\ &0\quad\text{otherwise}.\end{aligned}\right.$ (36)
  3. 2.
    Normalized Mutual Information (NMI) [42]: NMI is utilized to describe the mutual dependence between clustering results (after being mapped into the real labels) and ground truth labels. Given the clustering results $\mathbf{m}$ and the ground truth labels $\mathbf{l}$ (both including all samples), $I(\mathbf{m},\mathbf{l})$ denotes the mutual information between $\mathbf{m}$ and $\mathbf{l}$. $H(\mathbf{m})$ and $H(\mathbf{l})$ are the entropy of $\mathbf{m}$ and $\mathbf{l}$ respectively. The same as ACC, a larger NMI means better clustering performance.
    $\text{NMI}(\mathbf{m},\mathbf{l})=\frac{I(\mathbf{m},\mathbf{l})}{\sqrt{H(\mathbf{m})H(\mathbf{l})}}.$ (37)
  4. 3.
    Training time: For an algorithm, training time starts from the initialization and ends when the algorithm outputs the selected features.
For orientation-specific data such as multi-channel time series, the temporal dependency of the elements within the same channel makes the clustering metrics unreliable due to the feature independence hypothesis in K-means. Therefore, we define a metric as follows:
  1. [•]
  2. 1.
    Proportion of correctly selected features (POC): We compute POC as:
    $\text{POC}=c/h,$ (38)
    where $c$ is the number of correctly selected features, and $h$ is the total number of selected features at each execution. POC reflects a UFS method’s ability to select discriminative features. In the experiments, we select features with maximal between-class variance (BCV) as the correct ones. We compute POC on slice-wise data instead of ACC and NMI.
5.2 Datasets
Before detailing specific datasets, we first categorize tensor datasets based on their structural organization. As shown in Figure 2, third-order tensor datasets can be divided into two scenarios: tube-wise and slice-wise, according to their organization. Slice-wise data is orientation-specific and exhibits feature relations along the 1-mode, while tube-wise data exhibits feature relations along both the 1-mode and 2-mode.. This distinction provides a basis for evaluating the ability of UFS methods to utilize tensor structure information. In our evaluation, we compare clustering ACC and NMI on tube-wise data and POC on slice-wise data. Experimental results demonstrate the effectiveness of the proposed method in both scenarios.
Table 2: Two scenarios for third-order data tensors. Assume each frontal slice is a sample. Different colors refer to different variables. This paper regards variables as features. Tube-wise scenario: elements within one fiber belong to one variable. Slice-wise scenario: elements within one horizontal slice belong to one variable.
Scenario Visualization Example
Tube-wise [Uncaptioned image] 1. snapshots of array signal [8, 10, 11]. 2. images and video[7, 12, 13].
Slice-wise (orientation-specific) [Uncaptioned image] 1. time series with different channels [14, 15, 16, 9]. 2. multi-way statistical tables [17].
To preliminarily evaluate the UFS ability of our proposed method in the two tensor scenarios, we design two synthetic tensor datasets:
  1. [•]
  2. 1.
    Orbit: As an example of slice-wise data, Orbit comprises 3-dimensional orbit samples with two distinct radii. Each sample includes three channels representing orbit coordinates, with the remaining channels containing random Gaussian noise.
  3. 2.
    Array: As an example of tube-wise data, Array is constructed using a $10\times 10$ synthetic radar unit array that receives a series of echoes from a stationary target. Each echo is represented as a matrix of unit sampling values, treated as a sample. Random measurement errors are introduced to half of the echo matrices to create two distinct classes.
We evaluate our method on nine real-world datasets, including: 1) six tube-wise datasets: COIL20 [54], USPS [52], BreastMNIST [44], Imm40 [51], PIE [13] and JAFFE [43]; 2) four slice-wise datasets: UCIHAR [45] and UCIDSA [46], UMCS [^1] https://www.kaggle.com/c/seizure-detection/data and ASCH [53]. The detailed information of each dataset is shown in Table 3. PIE, Imm40, and JAFFE contain frontal facial images, labeled according to individuals. BreastMNIST is a medical dataset containing breast ultrasound images, from which we construct the data tensor using the test set. USPS is a dataset of handwritten digits. UCIHAR and UCIDSA are human activity datasets collected from multiple subjects using motion sensors. UCIDSA contains multi-channel signals, while UCIHAR comprises extracted statistical features (e.g. mean, std, max, entropy). UMCS and ASCH are electroencephalogram (EEG) datasets collected for disease research, capturing brain activity signals. Due to the large volume of chosen slice-wise data, we retrieve a subset of samples to construct datasets. For UCIDSA, we extract the first $125$ sample points from all $45$ channels covering $45$ activities. For UCIHAR, we select $9$ individuals with all $561$ features sampled during $6$ activities. For UMCS, we collect the first $125$ sample points in all $65$ channels of the first human patient, classified into ictal and interictal segments. For ASCH, we collect the first $640$ sample points in all $16$ channels from healthy groups and groups with symptoms. All datasets are normalized and bounded within $[-1,1]$ for consistency.
Table 3: Statistics of datasets.
Type Dataset #Feature #Samples #Classes
Tube-wise Array $10\times 10$ 100 2
USPS [52] $16\times 16$ 9298 10
BreastMNIST [44] $28\times 28$ 156 2
Imm40 [51] $32\times 32$ 240 40
COIL20 [54] $32\times 32$ 1440 20
PIE [13] $32\times 32$ 1166 53
JAFFE [43] $64\times 64$ 213 7
Slice-wise Orbit $9\times 41$ 100 2
UCIDSA [46] $45\times 125$ 152 19
UCIHAR [45] $561\times 9$ 273 6
UMCS 1 $65\times 125$ 174 2
ASCH [53] $16\times 640$ 84 2
5.3 Synthetic experiment
In this experiment, we conduct four typical UFS methods on synthetic datasets with the grid search strategy. The dataset formulation and the best experimental results of comparative methods are illustrated in Figure 3. On Orbit, we retrieve three top-ranked features in each method’s result and plot their samples in a subset of data. If a method selects the correct features (i.e. the three orbit coordinates), then the scatter plot will contain rings of two different radii. Otherwise, it will be difficult to distinguish the two classes. On Array, we retrieve the feature score map corresponding to the best POC for each method, analyzing the match degree between the top-ranked feature positions and the units with measurement error. We conclude the observations as follows.
  1. [•]
  2. 1.
    Tensor-based methods are generally more capable of analyzing the data structure than non-tensor-based methods. Both CPUFS and STPCA-MP select the orbit coordinates on Orbit. Although CPUFS fails to precisely locate the units with error on Array, its score map clearly shows a grid-like characteristic that connects the rows and columns to analyze feature relations, capturing a roughly correct attention region. A similar pattern also appears in STPCA-MP, where the column-wise relations are learned. In contrast, RNE and NLGMS isolate the features, losing the tensor structure information. Especially in RNE, the scores on Array seem to be randomly distributed.
  3. 2.
    Our proposed method performs the best in selecting discriminative features on the two synthetic datasets. STPCA-MP selects the correct features on both datasets. In particular, the score map on Array is relatively sparse in rows and columns. By taking advantage of data orientation, STPCA-MP can more easily mine the latent feature relations and locate discriminative features.
Figure 3: Synthetic experiments on Orbit and Array. The left part gives the composition of the two synthetic data tensors. The right part shows the feature selection results of comparative methods in a visualized manner.
5.4 Clustering experiment on tube-wise data
In this experiment, we perform clustering on five tube-wise real-world datasets. The regularization parameters of all methods go through a grid search in $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$. We set the selected features to $\{50,100,150,200,250,300\}$ on COIL20, JAFFE, BreastMNIST , Imm40, and $\{10,30,50,70,90,110\}$ on USPS. The best results and the corresponding training time of all methods are given in Table 4. The detailed ACC and NMI curves are illustrated in Figure 4. By analyzing the results, we come out with the following conclusions:
  1. [•]
  2. 1.
    Features selection improves data compactness. On each dataset, one or more methods are capable of achieving performance that approaches or even surpasses the baseline while utilizing only a subset of features. In particular, the proposed STPCA-MP achieves significantly higher clustering metrics than all features on Jaffe, BreastMNIST, and Imm40.
  3. 2.
    STPCA-MP delivers reliable and outstanding performance in clustering tasks. Firstly, compared to graph-based methods like SOGFS, STPCA-MP obtains more global data information by exploiting the reconstruction term, which guarantees the basic clustering performance. We can see the curves of STPCA-MP, on the whole, have a smaller oscillation in amplitude and share better consistency in trend between ACC and NMI. Besides, STPCA-MP reflects a more natural decreasing (or increasing) progression in variation pattern. Secondly, the T-SVDM-based sparse framework is beneficial for extracting tensor structure information, enabling STPCA-MP to describe the feature correlation in a more delicate manner, compared to CPD-based tensor methods (CPUFS and CPUFSnn), Tucker-based tensor method (MSPCA), and non-tensor-based methods (SPCAFS, RNE). Therefore, STPCA-MP demonstrates superior performance, consistently ranking at the forefront on all datasets.
  4. 3.
    STPCA-MP achieves the best performance at a smaller training time cost. On all datasets in Table 4, STPCA-MP takes less than $1$ second to complete training and feature selection, due to its slice-by-slice operation. Note that this is accomplished with reasonable prior knowledge of data organization. As a comparison, NLGMS, also showing remarkable average performance, brings much more computational cost as it overlooks the tensor structure by vectorization operation.
Table 4: Metrics of the comparative methods on real-world datasets. The best ACC, best NMI, and mean POC regarding grid-search range are shown. The training time corresponding to the best result is recorded in seconds. The best result on each dataset is boldfaced, while the second best one is underlined and italic.
Method PIE COIL20 JAFFE BreastMNIST Imm40 USPS UCIDSA UCIHAR UMCS ASCH
All Features ACC 26.21 ACC 58.34 ACC 20.16 ACC 58.25 ACC 48.43 ACC 65.29 POC - POC - POC - POC -
NMI 51.19 NMI 75.74 NMI 4.51 NMI 2.41 NMI 74.21 NMI 61.11
RNE[50] ACC 18.67 ACC 57.83 ACC 26.96 ACC 60.26 ACC 58.54 ACC 55.51 POC TIME 40.00 83.57 POC TIME 0 170.22 POC TIME 40.00 193.22 POC TIME 66.67 205.33
NMI 42.64 NMI 73.81 NMI 12.89 NMI 4.34 NMI 81.11 NMI 50.74
TIME 5.56 TIME 18.17 TIME 63.42 TIME 8.16 TIME 15.44 TIME 1.60
SOGFS[40] ACC 28.22 ACC 57.54 ACC 24.88 ACC 61.45 ACC 57.06 ACC 57.63 POC TIME 37.33 2409.95 POC TIME 1.60 5569.15 POC TIME 17.33 5274.80 POC TIME 23.33 8986.97
NMI 53.01 NMI 72.55 NMI 10.01 NMI 4.55 NMI 79.60 NMI 52.21
TIME 34.28 TIME 55.56 TIME 1794.41 TIME 37.12 TIME 83.77 TIME 391.60
SPCAFS[38] ACC 41.55 ACC 58.10 ACC 29.48 ACC 60.90 ACC 62.86 ACC 64.63 POC TIME 70.67 140.61 POC TIME 23.60 274.54 POC TIME 28.00 435.80 POC TIME 70.00 6569.55
NMI 65.04 NMI 74.02 NMI 12.07 NMI 5.14 NMI 84.09 NMI 61.13
TIME 18.37 TIME 3.11 TIME 59.36 TIME 0.61 TIME 15.76 TIME 0.34
NLGMS[56] ACC 38.40 ACC 61.30 ACC 30.94 ACC 61.86 ACC 59.51 ACC 66.08 POC TIME 40.53 222.98 POC TIME 17.08 410.78 POC TIME 21.87 782.96 POC TIME 46.00 3114.56
NMI 62.69 NMI 76.74 NMI 18.34 NMI 4.29 NMI 82.03 NMI 61.45
TIME 25.99 TIME 25.71 TIME 498.12 TIME 2.98 TIME 19.08 TIME 511.68
CPUFS[18] ACC 31.30 ACC 60.05 ACC 27.82 ACC 61.11 ACC 59.18 ACC 64.99 POC TIME 34.35 51.01 POC TIME 19.16 71.78 POC TIME 21.49 11.47 POC TIME 34.93 5.91
NMI 56.56 NMI 75.48 NMI 13.42 NMI 3.48 NMI 81.41 NMI 59.79
TIME 290.69 TIME 69.89 TIME 11.20 TIME 6.87 TIME 14.01 TIME 1559.53
CPUFSnn[18] ACC 32.51 ACC 61.47 ACC 26.15 ACC 61.54 ACC 57.85 ACC 62.79 POC TIME 34.40 48.77 POC TIME 23.41 17.79 POC TIME 23.75 12.66 POC TIME 34.27 6.67
NMI 32.51 NMI 75.85 NMI 11.86 NMI 4.03 NMI 79.18 NMI 58.37
TIME 101.76 TIME 20.72 TIME 11.68 TIME 7.11 TIME 13.35 TIME 485.50
MSPCA[33] ACC 30.66 ACC 59.74 ACC 27.51 ACC 63.53 ACC 62.54 ACC 63.37 POC TIME 86.67 0.37 POC TIME 24.20 3.22 POC TIME 30.80 0.02 POC TIME 69.00 0.15
NMI 53.81 NMI 75.84 NMI 10.47 NMI 3.46 NMI 83.40 NMI 59.86
TIME 0.33 TIME 0.02 TIME 0.01 TIME 0.01 TIME 0.07 TIME 0.05
STPCA-MP (ours) ACC 42.84 ACC 61.64 ACC 34.74 ACC 64.34 ACC 68.92 ACC 66.77 POC TIME 92.27 0.08 POC TIME 35.20 0.17 POC TIME 41.07 0.04 POC TIME 83.33 0.94
NMI 66.80 NMI 76.29 NMI 23.42 NMI 5.20 NMI 87.81 NMI 61.75
TIME 0.03 TIME 0.04 TIME 0.05 TIME 0.02 TIME 0.02 TIME 0.06
Figure 4: The ACC and NMI curves on tube-wise datasets. The upper row shows the ACC curves, while the bottom row displays the NMI curves.
5.5 POC comparison experiment on slice-wise data
For slice-wise datasets, we view each channel (row vector) as a feature. Therefore, we perform this experiment in the following steps: First, we implement a UFS algorithm and obtain the scores for each element. Second, we sum row-wise scores to evaluate features, using POC to measure the agreement between top-ranked features and those with maximal between-class variance (BCV). We compute BCV after the data are zero-centered.
The BCV of selected features on four slice-wise datasets is shown in Figure 5, where the results corresponding to the greatest POC of typical comparative methods are given. The mean POC of all methods is illustrated in Table 4. We have the following observations:
  1. [•]
  2. 1.
    For each dataset, there exists a small set of features with prominent BCV, judging by the peaks on the BCV curves in Figure 5. Greater BCV usually implies better inter-class discriminability.
  3. 2.
    SPCA-based methods are generally more capable of selecting the features of the greatest BCVs. SPCAFS, MSPCA, and STPCA-MP have a stabler performance in terms of POC, owing to the encoder-decoder framework that minimizes the reconstruction error as well as maximizes the variance among samples on the global manifold. Attributed to the sparse tensor decomposition, MSPCA also achieves relatively stable performance on slice-wise data, referring to its mean POC in Table 4.
  4. 3.
    STPCA-MP has the advantage of handling slice-wise data over comparative methods. Due to the exploitation of data orientation, STPCA-MP can better learn the intrinsic relationship between different channels in a slice-wise tensor, selecting a considerable amount of BCV peaks in Figure 5. Although the best result of NLGMS is compatible to our method on some datasets, STPCA-MP demonstrates enhanced stability in selecting discriminative features, according to the mean POC regarding grid-search range in Table 4. On average, STPCA-MP is more effective toward slice-wise data, compared to the state-of-the-arts.
Figure 5: The BCV of selected features on slice-wise datasets.
5.6 Visualization
In this experiment, we perform a visualization to empirically compare the interpretability of selected features. The best results of three tensor-based methods and one non-tensor-based method in terms of NMI on two image datasets: JAFFE and BreastMNIST, are visualized in Figure 6, where we highlight the selected features to see the attention region of the UFS methods. We have the following observations:
  1. [•]
  2. 1.
    STPCA-MP pays more attention to discriminative features. The exploitation of data orientation enables STPCA-MP to mine essential information with different spatial distributions. On JAFFE, STPCA-MP selects more horizontal features that cover the regions of eyes, eyebrows, mouth, and hair, etc. beneficial for distinguishing facial expressions, while CPUFS and MSPCA are misguided by background pixels. NLGMS, though it selects reasonable features as well, has a more scattered region and fails to select correlated features. On BreastMNIST, the features selected STPCA-MP tend to be grouped at an abnormal shadowed region composed of both horizontally and vertically distributed features, while the results of CPUFS, MSPCA and NLGMS are diverse. We can also observe that compared to CPUFS, MSPCA selects more rows (or columns) containing discriminative features and is less affected by the background, which we believe is attributed to the direct utilization of sparse factor matrices (CPUFS doesn’t apply a sparse tensor decomposition).
  3. 2.
    The score map of STPCA-MP exhibits significantly clearer and more distinct patterns than comparative methods. Thanks to the slice-by-slice operation, STPCA-MP describes feature correlations in a more delicate manner. On JAFFE, the score map of STPCA-MP approximates the shape of a human face. On BreastMNIST, the score map of STPCA-MP demonstrates a smooth score gradient transition radiating from the center of the shadowed region. In contrast, the score maps of CPUFS and MSPCA are less interpretable, due to the vague description of mode relations rooted in CPD and Tucker decomposition. As for NLGMS, its score map adds significant weights to the redundant features in the background. This phenomenon is possibly evoked by the sensitivity to noise features of the constructed graph.
Figure 6: Visualization on JAFFE and BreastMNIST. We show the score maps and highlight the positions of the top-ranked features in the samples. We set the number of selected features to $300$ on JAFFE, and $50$ on BreastMNIST.
5.7 Training time comparison
In this experiment, we compare the computational cost of comparative methods. The training time corresponding to the best clustering results is given in Table 4, while the average training time in terms of regularization parameter combinations is illustrated in Figure 7. We have the following observations:
  1. [•]
  2. 1.
    Tensor-based methods are generally more efficient on high-dimensional data. By leveraging tensor-based techniques, the computational complexity is reduced from quadratic or cubic in the full feature space to quadratic or cubic within a single mode dimension. Therefore CPUFS, CPUFSnn, MSPCA, and our proposed method averagely cost less training time on JAFFE, UMCS, and ASCH compared to other methods.
  3. 2.
    STPCA-MP is efficient on both high-dimensional and large-scale data. CPUFS and CPUFSnn usually take hundreds of iterations to converge because of the non-convexity. Moreover, their computational complexity scales quadratically with #$\#(Sample)$, whereas STPCA-MP exhibits linear complexity relative to #$\#(Sample)$. Therefore, STPCA-MP takes considerably less training time than CPUFS and CPUFSnn on all datasets, especially on large-scale datasets. Although MSPCA is on average more efficient than STPCA-MP due to a smaller weight parameter size requiring computation in Tucker decomposition, it gains no advantage over the proposed method in both clustering tasks and the POC metric. Moreover, the training time associated with optimal performance shows only a marginal discrepancy between MSPCA and STPCA-MP (See Table 4), which is in an acceptable range. On the whole, the proposed method is both effective and efficient.
(a) JAFFE (#Feature=$4096$, #Sample$=213$)
(b) UMCS (#Feature=$8125$, #Sample$=174$)
(c) ASCH (#Feature=$10240$, #Sample$=84$)
(d) USPS (#Feature=$784$, #Sample$=9298$)
Figure 7: Average training time during grid search.
5.8 Convergence analysis
In this experiment, we study the convergence of the proposed method. Figure 8 illustrates the converge curves of objective function value on four real-world datasets. Utilizing the HPSD projector, the proposed two-step algorithm converges fast into the solution space within dozens of iterations.
(a) PIE
(b) JAFFE
(c) BreastMNIST
(d) UCIDSA
Figure 8: Converge curves of STPCA-MP on JAFFE and BreastMNIST. All regularization parameters are set to 1.
5.9 Parameter sensitivity analysis
In this section, we analyze the impact of the parameters $\lambda$ and $\eta$ in STPCA-MP on the performance of feature selection. We vary the parameters in the range of $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$, set the selected features to 100, and record the ACC and NMI of STPCA-MP with all possible parameter combinations. The results on JAFFE and USPS is illustrated in Figure 9. We can see that
  1. [•]
  2. 1.
    On JAFFE, the performance of STPCA-MP is relatively stable when $\lambda$ and $\eta$ vary in a certain range, and degenerates when the parameters become too large. This tendency is consistent between ACC and NMI.
  3. 2.
    On USPS, STPCA-MP is generally not sensitive to the choice of parameters, achieving relatively stable performance with respect to both ACC and NMI.
(a) JAFFE (ACC)
(b) JAFFE (NMI)
(c) USPS (ACC)
(d) USPS (NMI)
Figure 9: Parameter sensitivity analysis of STPCA-MP on JAFFE and USPS.
5.10 Discussion on different choices on data orientation and transform domain.
In this experiment, we discuss the influence of data tensor orientation and transform domain. We conduct STPCA-MP with different data tensor rotations and transform matrices $\mathbf{M}$ on several datasets. The detailed ACC curves on JAFFE and USPS are illustrated in Figure 10. We further provide the best ACC and the mean POC in Table 5. The visualized result with $\mathbf{M}=\mathbf{I}$ on PIE and JAFFE is given in Figure 11. We have the following observations:
  1. [•]
  2. 1.
    Different data orientations can significantly affect STPCA-MP’s clustering performance. We can see significant gaps among certain ACC curves in Figure 10. Like looking at an identical object from different views, decomposing a data tensor in different directions naturally leads to discrepant information. The visualized result in Figure 11 implies that STPCA-MP can capture features with different spatial distributions given different data orientations. A proper choice of data orientation is beneficial for mining feature correlations. The spatial distribution and latent correlation structure of discriminative features are heterogeneous among data tensors. On UCIDSA and UMCS, the overall performance of the Dir1s is better than the Dir2s, since the feature relations in multi-channel time-series data exist along 1-mode. In conclusion, performing slice-by-slice operations along a certain direction benefits UFS at multiple levels on different datasets.
  3. 2.
    A suitable transform domain can benefit the performance. Since the transform extracts the information from one of the two feature modes, while a slice-by-slice operation is conducted along the other feature mode, a proper transform domain can integrate more feature correlations as prior knowledge for the following optimization and UFS. In Table 5, Dir1-EigVector sometimes achieves better results than Dir-I, especially on time-series data. However, there is no absolute conclusion that certain settings of transform domain necessarily outperform others. As we can observe in both Figure 10 and Table 5, there is no setting under which the metrics consistently outperform all other configurations. Identical settings of orientation and transform domain can lead to substantial performance discrepancies across different datasets. In general, the choice of orientation and transform domains requires detailed configurations to achieve the best performance.
(a) JAFFE (different orientations)
(b) USPS (different orientations)
(c) JAFFE (different $\mathbf{M}$)
(d) USPS (different $\mathbf{M}$)
Figure 10: The clustering ACC of STPCA-MP with different data tensor orientations and $\mathbf{M}$. Different data tensor orientations. ’Dir1’ means $D^{o}=\{1,3,2\}$. ’Dir2’ means $D^{o}=\{2,3,1\}$. ’I’ means $\mathbf{M}=\mathbf{I}$. ’DFT’ means that $\mathbf{M}$ is the discrete Fourier matrix (we conduct fast Fourier transform in practice). ’EigVector’ means $\mathbf{M}$ is the eigenvectors decomposed from the self-correlation matrix of the other feature mode. ’Random’ means $\mathbf{M}$ is a random orthogonal matrix. (c) and (d) are obtained by running STPCA-MP-Dir1.
Figure 11: Visualized feature selection results of our proposed method with different data orientations. The experimental condition is the same as in Figure 6.
Table 5: The best ACC and mean POC of our proposed method with different data orientations and transform domains. The best result on each dataset is boldfaced, while the second best one is underlined and italic.
Setting Metric JAFFE BreastMNIST USPS PIE Metric UCIDSA UMCS
Dir1-I ACC 35.29 64.34 63.32 42.84 POC 92.27 41.07
Dir1-DFT ACC 31.35 55.13 64.13 28.20 POC 9.33 38.67
Dir1-EigVector ACC 30.49 57.56 63.7 30.98 POC 92.8 44.80
Dir2-I ACC 30.74 62.44 66.77 36.87 POC 65.07 28.8
Dir2-DFT ACC 28.48 62.29 57.34 38.39 POC 24 16.53
Dir2-EigVector ACC 22.68 65.32 61.55 33.02 POC 45.87 18.13
6 Conclusion
In this paper, we propose a novel Sparse Tensor PCA method named STPCA-MP by combining the encoder-decoder structure of SPCA and the orientation-dependent framework of T-SVDM. We deduce a convex version of STPCA-MP that utilizes the sparse reconstruction tensor to perform UFS. An efficient slice-by-slice algorithm to conduct STPCA-MP is proposed to guarantee fast convergence. The experimental results on real-world data show that our method outperforms state-of-the-art methods on average and can handle different data organization scenarios at a smaller training time cost. In particular, STPCA-MP has the advantage of selecting discriminative features with the greatest BCV on slice-wise data. In addition, we also test the parameter sensitivity, examine the efficient convergence, and discuss the impacts of choosing different data orientations and transform domains. The experimental results show that with a suitable combination of data orientation and transform domain, our method can capture abundant discriminative information and achieve remarkable performance.
There are two valuable research points concerning STPCA-MP in the future. First, the best choice of transform domain is possibly related to the background knowledge of data. More experiments should be carried out for further applications. Second, the extension to higher-order scenarios of STPCA-MP still lacks an effective and efficient optimization algorithm. We will be working on addressing the above two problems in the future.
Appendix A Appendix
[…]
pf3Proof of Theorem 1
{pf3}
To simplify the deduction, we first rewrite the $i$-th problem with compact symbols:
$\min\limits_{\mathbf{U},\mathbf{Q}}$ ${\|\mathbf{X}-\mathbf{U}\mathbf{Q}^{H}\mathbf{X}\|}^{2}_{F}+\lambda{\|\mathbf{Q}\|}_{2,1},$ (39)
$s.t.$ $\mathbf{U}^{H}\mathbf{U}=\mathbf{I}_{k}.$
Thus, we can rewrite the objection function as:
${\rm Tr}\left(\mathbf{X}\mathbf{X}^{H}\right)-\sum\limits^{k}_{j=1}\bigg{[}2Tr\left(\mathbf{u}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}\right)-\frac{\lambda}{\|\mathbf{q}_{j}\|_{2}}\mathbf{q}_{j}^{H}\mathbf{q}_{j}-{\rm Tr}\left(\mathbf{q}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}\right)\bigg{]}$ (40)
$={\rm Tr}\left(\mathbf{X}\mathbf{X}^{H}\right)-\sum\limits^{k}_{j=1}\bigg{(}2\mathbf{u}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}-\frac{\lambda}{\|\mathbf{q}_{j}\|_{2}}\mathbf{q}_{j}^{H}\mathbf{q}_{j}-\mathbf{q}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}\bigg{)}$
if we view the above problem as a sum of $k$ subproblems with respect to $\mathbf{u}_{j}$ and $\mathbf{q}_{j}$. Then in each iteration, given a fixed $\mathbf{u}_{j}$, we can have each subproblem minimized at
$\mathbf{q}^{*}_{opt}=\left(\mathbf{X}^{*}_{L_{e}}\mathbf{X}^{T}_{L_{e}}+\frac{\lambda}{\|\mathbf{q}_{j}\|_{2}}\mathbf{I}_{q}\right)^{-1}\mathbf{X}^{*}\mathbf{X}^{T}\mathbf{u}_{j}^{*}.$ (41)
By substituting (41) back to (39) we can obtain
$\mathbf{u}_{opt}=$ (42)
${\underset{\mathbf{u}^{H}\mathbf{u}=1}{\arg\min}}-\mathbf{u}\mathbf{X}\mathbf{X}^{H}\left(\mathbf{X}\mathbf{X}^{H}+\frac{\lambda}{\|\mathbf{q}\|_{2}}\mathbf{I}_{q}\right)^{-1}\mathbf{X}\mathbf{X}^{H}\mathbf{u}^{H}.$
Given the singular decomposition $\mathbf{X}=\mathbf{V}\mathbf{D}\mathbf{M}^{H}$, we have $\mathbf{u}^{opt}_{j}=s_{j}\mathbf{v}_{j}$ where $s_{j}=1$ or $-1$. Then, we obtain $\mathbf{q}^{opt}_{j}=s_{j}d_{jj}/(\|d_{jj}\|^{2}_{2}+{\lambda}/{\|\mathbf{q}_{j}\|_{2}})\mathbf{v}_{j}$.
Finally, we have
$\mathbf{U}_{opt}\mathbf{Q}^{H}_{opt}$ $=\sum\limits^{k}_{j=1}\mathbf{u}_{j}^{opt}(\mathbf{q}_{j}^{opt})^{H}=\mathbf{V}\mathbf{\Sigma}\mathbf{V}^{H}\in H^{q}_{+},$ (43)
[…]
Eventually, after each subproblem of (15) is solved, the optimal solution satisfies $\hat{\mathcal{A}}^{opt}_{::i}=\hat{\mathcal{U}}^{opt}_{::i}(\hat{\mathcal{Q}}^{opt}_{::i})^{H}\in H^{q}_{+}$.
[…]